A merchant blends tea that sells for $4.95 a pound with tea that sells for $2.55 a pound to produce 60 lb of a mixture that sells for $3.75 a pound. How many pounds of each type of tea does the merchant use in the blend?
Answer by mananth(16946)

Price ---------------- quantity
Tea type I 4.95 ---------------- x pounds
Tea type II 2.55 ------ 60 - x pounds
Mixture 3.75 ---------------- 60
60
4.95 x + 2.55 ( 60 - x ) = 225
4.95 x + 153 - 2.55 x = 225
4.95 x - 2.55 x = 225 - -153
2.4 x = 72
/ 2.4
x = 30 pounds 4.95% Tea type I
30 pounds 2.55% Tea type II

m.ananth@hotmail.ca